OpenStudy (anonymous):

I need help solving this definite integral, from x = -2 to x=e: f(x)=x-e, if x>0 and f(x)=e^(x+2), if x<=0 Thanks !

OpenStudy (turingtest):

you need to split this up into two integrals where are you having trouble?

OpenStudy (anonymous):

I got final answer being -2+e^(e+3). I tried to check it but it was wrong according to geogebra :S

OpenStudy (turingtest):

is that first function\(f(x)=x-e\) or \(f(x)=x^{-e}\) ?

OpenStudy (anonymous):

\[f(X)=x-e, if x>0 \\ f(x)=e^{x+2}, if x\le 0\]

OpenStudy (turingtest):

so what did you put for the first integral?

OpenStudy (anonymous):

we split into integral fropm -2 to near 0 (x-e) and from 0 to e (e^(x+2))

OpenStudy (turingtest):

good, so\[\int_{-2}^0x-edx+\int_0^ee^{x+2}dx\]right?

OpenStudy (anonymous):

exactly

OpenStudy (turingtest):

\[\int_{-2}^0x-edx+\int_0^ee^{x+2}dx=\frac{x^2}2-ex|_{-2}^0+e^{x+2}|_0^e=2+2e+e^{e+2}-e^2\]

OpenStudy (anonymous):

I tried to check, it is not checking, doing the calculation in geogebra, or wolfram alpha